Thomas’ Calculus Early Transcendentals 13ed

Study guide for quiz 06

Quiz problems include both the lecture contents and homework problems.

1. Section 4.4: Review procedures for sketching a curve. Such as how to determine whether the curve is increasing/decreasing and concave up/down.

2. Section 4.5: Review how to determine a indeterminate form using L’Hôpital’s Rule. Go through all examples in section 4.5. Find one example to each of the form ∞ − ∞, 0 · ∞, 1^{∞}, 0^{0}, ∞^{0} and study how to find the limit by L’Hôpital’s Rule.

3. Section 4.5: Find an example of the form \frac{0}{0} that L’Hôpital’s Rule does not leads to an

answer, but after rewriting \frac{0}{0} = \frac{\frac{1}{0}}{\frac{1}{0}} = \frac{∞}{∞}, L’Hôpital’s Rule can be applied to find the limit.